A sinusoidal wave moving along a string is shown twice below, as crest A travels in the positive direction of an x axis by distance d = 1.5 cm in 4.4 ms. The tick marks along the axis are separated by 3.0 cm; height

H = 6.00 mm.

The wave equation is of the form below.

y(x, t) = ym sin(kx ± ωt)

(a) What is ym?
mm

(b) What is k?
rad/m

(c) What is ω?
rad/s

(d) What is the correct choice of sign in front of ω?

+−    
Answer #1

d = 1.5 cm = 0.015 m

t = 4.4 ms = 4.4\times 10^{-3} s

x = 3.0 cm = 0.03 m

H = 6.0 mm = 6.0\times 10^{-3} m

wave speed is v = \frac{d}{t} = \frac{0.015 m}{4.4\times 10^{-3}s}

v =3.409 \frac{m}{s}

(a)

y_{m} = \frac{H}{2} = \frac{6.0 mm}{2}

y_{m} = \frac{6.0 mm}{2}

y_{m} =3.0 mm

(b)

wave length \lambda = 4x

\lambda = 4\times 0.03 m

\lambda = 0.12 m

k = \frac{2\pi }{\lambda}

k = \frac{2\pi }{0.12 m}

k = 16.667\pi \frac{rad}{m}

k = 52.36\frac{rad}{m}

(c)

\omega = vk

\omega = 3.409 \frac{m}{s}\times 52.36 \frac{rad}{s}

\omega = 178.5 \frac{rad}{s}

(D)

the wave move +x direction, the wave equation is y = y_{m}sin\left ( kx-\omega t \right )

Ans ===> (-)